Stan paints twice as fast as Daniel. Working together, Stan and Daniel can paint 2,400 square feet in 4 hours. Another employee, Roger, joined their painting and together, Stan, Daniel & Sue can paint 3,600 square feet in 3 hours.
If Roger works alone, how many square feet can he paint in 4 hours and 27 minutes?

Answer by richwmiller(17219)

s=2d
s+d=3d
3d=2400 in 4 hours
d=800 in 4 hrs 200 in 1 hr
s=1600 in 4 hrs 400 in one hour
3*200+3*400=600+1200=1800 for s and d in 3 hours
which means that Roger and Sue can do 1800 sq ft in 3 hours.
We don't have enough info to know how Roger works alone.

If you meant roger and sue alone then ask again.